OpenStudy (anonymous):

Evaluate (7/16)^-2 A. 49/245 B. -49/256 C. 256/49

OpenStudy (anonymous):

C

OpenStudy (blurbendy):

It's C because you can rewrite (7/16)^(-2) as: 1 / (7/16)^2 = 1 / (49/256) = 256 / 49